### v1

@@ 5 Partial Exclusion and Monotonicity
(Partial Exclusion Restriction) For $d\in\{0,1\}$, $Y_{d}(z)=Y_{d}(z_{c})$ for all $z\in\mathcal{Z}$.
This partial exclusion assumption relaxes Assumption 2.2. It allows the potential outcomes to depend on judges’ observable characteristics but not their fixed effects or identity. For instance, when the treatment of interest is incarceration, judges could assign and differ in other punishments like probation, fines, or sentence length. These other punishments could directly affect potential outcomes, making Assumption 2.2 unlikely. For example, minority judges may be less lenient in their sentence length than their majority counterparts (Johnson, 2014). Beyond the decision to incarcerate, different sentence lengths may have divergent effects on, for example, later labor market outcomes. If the sentence length is not observed or controlled, we would expect the potential outcome to depend on whether a judge is a minority judge through this channel. The partial exclusion assumption says whether or how the judge assigns other types of punishment only depends on the judge’s observable ($Z_{c}$), but not the judge’s identity ($Z_{I}$). In other words, a defendant will end up with the same pair of potential outcomes $(Y_{1}(Z_{c}),Y_{0}(Z_{c}))$ as long as he or she is assigned to judges with the same observed $Z_{c}$. Finally, when the only instrument variable we observe in the data is the identity of the judge $Z_{I}$, then the partial exclusion assumption is equivalent to the original exclusion Assumption 2.2. The next assumption relaxes the monotonicity condition.
@@
Mogstad, Torgovitsky, and Walters (2019) introduced partial monotonicity assumption, which significantly weakens Assumption 2.3 since it does not require comparing the level of leniency across judges with different observable characteristics. For instance, let $Z_{c}=(Z^{R}_{c},Z^{P}_{c})$ be composed of the following binary variables: $Z^{R}_{c}$ equal to $1$ if the judge is black or Hispanic and 0 if not, while $Z^{P}_{c}$ is $1$ if the judge is from the Republican party and 00 if from the Democratic party. Imposing Assumption 2.3 means it is not possible to have a black democrat judge be more lenient than a white republican judge for some defendants while being less lenient for other defendants, i.e. these two judges may have different cut-off points, but they rank all the defendants by the same order. Mathmatically, we can not have both $\mathbb{P}(1=D(z_{I},1,0)>D(z_{I}^{\prime},0,1)=0)>0$ and $\mathbb{P}(1=D(z_{I}^{\prime},0,1)>D(z_{I},1,0)=0)>0$. However, there is a large empirical evidence of heterogeneity in the ranking of judges’ leniency across different types of offence or defendants (see Abrams, Bertrand, and Mullainathan, 2012; Stevenson, 2018). This is, however, compatible with the partial monotonicity. Its main advantage is it no longer requires a uniform ranking of defendants across different judges. Judges’ rankings are allowed to vary with their characteristics $Z_{c}$.
@@
Type $Z=(Z_{I},Z_{c})$’s judge treatment assignment mechanism is governed by the following threshold crossing model $D=1\{\nu(Z_{I},Z_{c})\geq U_{Z_{c}}\}$ for a measurable function $\nu$, where the distribution of $U_{z_{c}}$ is absolutely continuous for all $z_{c}\in\mathcal{Z}_{c}$.
Under Assumptions 2.1 and 5.3, we can rewrite the partial STC without loss of generality as follows:
$D(z_{I},z_{c})=1\left\{F_{U_{z_{c}}}(\nu(z_{I},z_{c}))\geq F_{U_{z_{c}}}(U_{z_{c}})\right\}\equiv 1\left\{P(z_{I},z_{c})\geq V_{z_{c}}\right\},$
@@
The key difference between the STC and the Partial STC is even though $V_{z_{c}}$ follows $Uniform[0,1]$ distribution, each defendant does not have a single $V$. Instead, he or she faces a collection of $\{V_{z_{c}},z_{c}\in\mathcal{Z}_{c}\}$. This unobserved latent variable is now different for judges with distinct observable characteristics. The partial STC has a natural interpretation as an extension of the Roy model (Canay, Mogstad, and Mountjoy, 2020). We can interpret $P(z_{I},z_{c})$ as the perceived gain of incarcerating a defendant by a type $z=(z_{I},z_{c})$ judge, and $V_{z_{c}}$ as the expected cost (but unobserved to the econometrician) of an incarcerating a defendant. The particularity of the partial STC is the expected cost can vary across judges with distinct observable characteristics $z_{c}$ but fixed within judges with the same $z_{c}$. In standard monotonicity assumption, the cost $V$ would be the same regardless of the characteristics $(z_{I},z_{c})$.
The partial STC has an alternative interpretation, allowing decision-makers (judges) to differ in their preferences and skills. Chan, Gentzkow, and Yu (2022) argue the standard STC (Assumption 2.4) implies the data must be consistent with all judges having the same signal $U$ (the same skill). However, under the partial STC, judges with distinct $Z_{c}$ are allowed to differ in their signal $U_{z_{c}}$ and thus, in their skill. In consequence, Assumption 5.3 significantly enhances the pattern of heterogeneity between judges in the judge leniency design.
Let us consider a situation where eight judges decide whether to incarcerate a given defendant to further elucidate the richer heterogeneity enabled by the partial monotonicity (or equivalently, the partial STC) assumption. We consider the two observable characteristics of the judges introduced earlier, $Z_{c}\equiv(Z_{c}^{R},Z_{c}^{P})\in\{0,1\}\times\{0,1\}$. These two binary observable characteristics lead to four judges’ types. The eight judges are evenly allocated across these four types.
@@
The left rectangle of Figure 6 shows the benefit and the expected cost of incarcerating the defendant on a separate unit segment for each judge. For example, $p_{11}$ and $p_{11}^{\prime}$ are the benefit of the two black democratic judges with type $Z_{C}=(1,1)$ to incarcerate the defendant. The right rectangle of Figure 6 plots the benefit numbers of all eight judges on the same unit segment. Similarly, $U_{11}$ represents the expected cost of incarcerating the defendant by a black democratic judge: they share the same expected cost or skills. A judge incarcerates the defendant when the corresponding benefit is higher than the expected cost of incarceration. In Figure 6, the judges who incarcerate the defendant are blue-coloured, while those of judges who release the defendant are red-coloured.
The behaviour of the eight judges does not violate Assumption 5.2 or Assumption 5.3. However, the standard monotonicity Assumption 2.3 is clearly violated (right rectangle of Figure 6). Indeed, the judge with benefit score $p_{11}^{\prime}$ incarcerates the defendant (blue-coloured), whereas judges with higher benefit scores $p_{01}$, $p_{10}$, or $p_{00}$ do not incarcerate the defendant (red-coloured). Note that Assumption 2.3 would not be violated for this group of judges only under one of these two conditions: (i) all four $V_{z_{c}}$ are greater than the maximum of the eight propensities or smaller than the minimum of all eight properties. In other words, when all judges make the same decision regarding this defendant, or (ii) judges who do not incarcerate the defendant must have lower benefit scores than judges who incarcerate the defendant. Moreover, one of these two conditions must hold for all defendants when we impose Assumption 2.3 (or Assumption 2.4).
@@
Assumptions 2.1, 5.1 and 5.3 are weaker than Assumptions 2.1, 2.2 and 2.3. We show that under these weaker conditions, it is still possible to identify meaningful treatment effect parameters.
Theorem 5.3 (Identification of causal effects under Partial exclusion and monotonicity)
If Assumptions 2.1, 5.1 and 5.3 hold, then:
@@
    (Identification of the LATE). Let $\mathcal{P}_{z_{c}}$ be the support of $P(Z_{I},Z_{c})$ conditioning on $Z_{c}=z_{c}$. Then for any pair $(p_{z_{c}},p_{z_{c}}^{\prime})\in\mathcal{P}_{z_{c}}\times\mathcal{P}_{z_{c}}$ such that $p_{z_{c}}<p_{z_{c}}^{\prime}$ we have the following identification results:
    $\frac{\mathbb{E}[g(Y)|P=p_{z_{c}}^{\prime},Z_{c}=z_{c}]-\mathbb{E}[g(Y)|P=p_{z_{c}},Z_{c}=z_{c}]}{p_{z_{c}}^{\prime}-p_{z_{c}}}\\ =\mathbb{E}[g(Y_{1}(z_{c}))-g(Y_{0}(z_{c}))|1\{p_{z_{c}}<V_{z_{c}}\leq p^{\prime}_{z_{c}}\}].$
@@
    (Identification of the MTE). For any $p_{z_{c}}\in\mathcal{P}_{Z_{c}}$ such that $\mathbb{E}[g(Y)|P=\cdot,Z_{c}=z_{c}]$ is continuously differentiable in the neighborhood of $p_{z_{c}}$, then,
    $\frac{\partial\mathbb{E}[g(Y)|P=t,Z_{c}=z_{c}]}{\partial t}\big{|}_{t=p_{z_{c}}}=\mathbb{E}[g(Y_{1}(z_{c}))-g(Y_{0}(z_{c}))|V_{z_{c}}=p_{z_{c}}].$
@@
    (Testable restrictions). For any fixed $z_{c}\in\mathcal{Z}_{c}$, $\mathbb{P}(y<Y\leq y^{\prime},D=1|P=p,Z_{c}=z_{c})$ and $-\mathbb{P}(y<Y\leq y^{\prime},D=0|P=p,Z_{c}=z_{c})$ are non-decreasing in $p$ for all $p\in\mathcal{P}_{Z_{c}}$ and any $y,y^{\prime}\in\mathcal{Y}$.
The proof of Theorem 5.3 is similar to Theorem 2.1 after conditioning on $Z_{c}=z_{c}$ and therefore omitted. The identification results stated in Theorem 5.3 (i)-(ii) demonstrate whenever there are two judges with distinct $Z_{I}$ but share the same observed characteristics $Z_{c}=z_{c}$, the conditional Wald estimand identifies the LATE provided the propensity scores for these two judges are different. Moreover, when the distribution of $Z_{I}|Z_{c}=z_{c}$ allows one to take the derivative of $\mathbb{E}[g(Y)|P=\cdot,Z_{c}=z_{c}]$, the conditional LIV estimand identifies the MTE. This identification result is a local version of the standard LATE and MTE identification.
@@
which is a total of fourteen inequalities. However, when invoking our weaker set of assumptions, we have only eight inequalities which characterize the sharp testable implications:
@@
The comparison of the testable implications in Theorems 2.1 and 5.3 confirms the judge leniency design is more stringent than the conditional judge leniency design. Hence, whenever the standard judge leniency design is rejected, the researcher may rely on its relaxed versions as long as the testable implications derived in Theorem 5.3 are satisfied.

### v2

@@ 5.1 Average Exclusion and Monotonicity
Let $\lambda_{z}=\Pr(Z=z)$, $p_{z}=\mathbb{E}[D_{z}]$, $p=\sum_{z\in\mathcal{Z}}\lambda_{z}p_{z}$, $\bar{D}=\sum_{z\in\mathcal{Z}}\lambda_{z}D_{z}$, and $\bar{Y}_{d}=\sum_{z\in\mathcal{Z}}\lambda_{z}Y_{d}(z)$ for $d\in\{0,1\}$.
  1. (a)
    Average exclusion restriction: E [ _z Z λ_z ( p_z - p ) { ( Y_0(z) - ¯Y_0 ) (1-D_z) + ( Y_1(z) - ¯Y_1 ) D_z } ] =0.
  2. (b)
    Average monotonicity: $\omega\equiv\sum_{z\in\mathcal{Z}}\lambda_{z}\left(p_{z}-p\right)\left(D_{z}-\bar{D}\right)\geq 0$ almost surely.
Under Assumptions 2.1 and 5.1, FLL’s Theorem 3 shows that the 2SLS estimand (of using $P(Z)$ as IV) has a causal interpretation since it can be written as a weighted average of the following treatment effect $\delta=\bar{Y}_{1}-\bar{Y}_{0}$ i.e.,
$\frac{\text{Cov}(Y,P(Z))}{\text{Cov}(D,P(Z))}=\mathbb{E}\left[\frac{\omega}{\mathbb{E}[\omega]}\delta\right].$ (5.1)
Note that the $\delta$ in Equation 5.1 is a deterministic function of the collection of potential outcomes $\{Y_{d}(z)\}_{d=0,1,z\in\mathcal{Z}}$. The 2SLS estimand is causal because it is a weighted average of $\delta$ and the weight $\omega$ is positive by the average monotonicity (Assumption 5.1-(b)).
Proposition 5.1 below provides a generalization and more transparent discussion of the FLL’s Theorem 3. First, we demonstrate that the average exclusion assumption is essentially equivalent to a zero covariance condition. Second, we show that Equation 5.1 indeed holds for any deterministic function of $\{Y_{d}(z)\}_{d=0,1,z\in\mathcal{Z}}$, not just for $\delta$.
To clarify these points, let us define $\alpha_{z}\equiv Y_{1}(z)-Y_{0}(z)$, and $\tilde{\alpha}=\sum_{z\in\mathcal{Z}}\alpha_{z}1\{Z=z\}=Y_{1}(Z)-Y_{0}(Z)$. Let $\alpha\equiv h(Y_{1}(1),...,Y_{1}(J),Y_{0}(1),...,Y_{0}(J))$ be an arbitrary measurable deterministic function of the collection of potential outcomes. $\delta$ defined in Assumption 5.1 is a special case when we pick $h(Y_{1}(1),...,Y_{1}(J),Y_{0}(1),...,Y_{0}(J))=\bar{Y}_{1}-\bar{Y}_{0}$. One could instead be interested in different treatment effects specific to each judge: $\alpha=\alpha_{z}$, $z=1,2,\cdots,J$. $\alpha$ can also be a quantity without clear economic interpretation such as $\alpha=\sum_{z\in\mathcal{Z}}zY_{1}(z)$.
Proposition 5.1
  1. (a)
    Under Assumption 2.1, and Assumption 5.1(b), the following equation holds for any measurable deterministic function $\alpha$ of the collection of potential outcomes:
    $\frac{\text{Cov}(Y,P(Z))}{\text{Cov}(D,P(Z))}=\mathbb{E}\left[\frac{\omega}{\mathbb{E}[\omega]}\alpha\right]+\frac{\text{Cov}\left((\tilde{\alpha}-\alpha)D+\tilde{Y}_{0},P(Z)\right)}{\mathbb{E}\left[\omega\right]}$
    where $\tilde{Y}_{d}=\sum_{z\in\mathcal{Z}}Y_{d}(z)1\{Z=z\}$.
  2. (b)
    Under Assumption 2.1, for $\alpha=\bar{Y}_{1}-\bar{Y}_{0}=\sum_{z}\lambda_{z}(Y_{1}(z)-Y_{0}(z))=\sum_{z}\lambda_{z}\alpha_{z}$ we have: Cov( (~α - α) D + ~Y_0,P(Z))=E [ _z Z λ_z ( p_z - p ) { ( Y_0(z) - ¯Y_0 ) (1-D_z) + ( Y_1(z) - ¯Y_1 ) D_z } ].
The proof for the proposition is collected in Section B.3. Under Assumption 2.1 (independence), Proposition 5.1(b) shows that the average exclusion restriction of FLL is, indeed, a special case of the zero-covariance assumption when $\alpha=\delta$. Proposition 5.1(a) further shows that if one targets an arbitrary quantity $\alpha\equiv h(Y_{1}(1),...,Y_{1}(J),Y_{0}(1),...,Y_{0}(J))$, and if one is willing to impose the same zero-covariance assumption on $\alpha$:
$\text{Cov}\left((\tilde{\alpha}-\alpha)D+\tilde{Y}_{0},P(Z)\right)=0,$ (5.2)
then one can always interpret 2SLS estimand as the weighted average of $\alpha$ with positive weights under the average monotonicity Assumption 5.1(b). This happens because the above zero-covariance condition in Equation 5.2 is a reduced-form condition, which assumes that the correlation between a reduced-form error (involving the parameter of interest) and the propensity score, i.e. $\text{Cov}\left(Y-\alpha D,P(Z)\right)=0$.
How does one assess the plausibility of the average exclusion condition? FLL provides a heuristic argument. [^19] Frandsen, Lefgren, and Leslie (2023, page 19): “Average exclusion can be probed by examining the correlation between judge-level treatment propensity and judge-level averages of alternative channels through which judges may affect outcomes if such channels are observed. Average exclusion may be more plausible if these correlations are near zero. However, this argument could also be invoked by anyone who wants to impose that Equation 5.2 holds for other $\alpha\neq\delta$. Also, it is difficult to justify why $\text{Cov}\left((\tilde{\delta}-\delta)D+\tilde{Y}_{0},P(Z)\right)=0$ but $\text{Cov}\left((\tilde{\alpha}-\alpha)D+\tilde{Y}_{0},P(Z)\right)\neq 0$ for other $\alpha\neq\delta$.
Furthermore, it is worth noting that the average exclusion restriction is not invariant to a relabelling of the treatment. In other terms, this assumption may hold if the researcher defines the treatment as $D$ equals $1$ if incarceration and 0 if not, while it may not hold if the researcher recodes the treatment as $D$ equals $1$ if no incarceration and 0 if incarceration. Indeed, after a relabelling, the zero-covariance in Equation 5.2 becomes: $\text{Cov}\left((\tilde{\alpha}-\alpha)D+\tilde{Y}_{1},P(Z)\right)=0$. It follows that the average exclusion assumption is invariant to a relabelling if and only if $\text{Cov}\left(\tilde{Y}_{1},P(Z)\right)=\text{Cov}\left(\tilde{Y}_{0},P(Z)\right)$.
Despite all those discussed above, we do observe a direct way to assess the validity of Assumptions 2.1 and 5.1. In fact, under these assumptions, there is:
$\left|\frac{\text{Cov}(Y,P(Z))}{\text{Cov}(D,P(Z))}\right|\leq\mathbb{E}\left[\left|\frac{\omega}{\mathbb{E}[\omega]}\delta\right|\right]\leq\mathbb{E}\left[\left|\delta\right|\right]\leq U-L$
where $U$ and $L$ are, respectively, the upper and lower bounds of $\mathcal{Y}$. Therefore, if the support of the outcome is bounded from both above and below, then the absolute value of the 2SLS estimand must also be bounded.
5.2 Conditioning on Judge’s Characteristics
In practice, it is not uncommon for researchers to have good reason to believe the assumptions hold after controlling for the judge’s specific characteristics. We explore this idea in this section and demonstrate that it is closely related to the partial exclusion assumption (defined below) and the partial monotonicity assumption made in Mogstad, Torgovitsky, and Walters (2019). Specifically, we decompose $Z$ into two components: $Z_{I}$ and $Z_{c}$, and we assume the monotonicity and exclusion restriction hold conditionally on $Z_{c}$. Here, $Z_{c}$ can be a judge’s race or political party, and $Z_{I}$ is a vector of the remaining characteristics.
Assumption 5.2 (Partial Exclusion)
Let $Z\equiv(Z_{I}^{\prime},Z_{c}^{\prime})^{\prime}$. For $d\in\{0,1\}$, $Y_{d}(z)=Y_{d}(z_{c})$ for all $z\in\mathcal{Z}$.
@@ 5.2 Conditioning on Judge’s Characteristics
The partial exclusion assumption relaxes Assumption 2.2 and allows the potential outcomes to depend on the subvector $Z_{c}$. For instance, when the treatment of interest is incarceration, judges could assign and differ in other punishments, such as probation, fines, or sentence length. These other punishments could directly affect potential outcomes, making Assumption 2.2 unlikely. Minority judges may be less lenient in their sentence length than their majority counterparts (Johnson, 2014). Beyond the decision to incarcerate, different sentence lengths may have divergent effects on future labor market outcomes. If the sentence length is not observed or controlled, we would expect the potential outcome to depend on whether a judge is a minority judge through this channel. The partial exclusion assumption states that whether and how the judge assigns other types of punishment depends only on a subset of the judge’s observable characteristics ($Z_{c}$), but not on others ($Z_{I}$). In other words, a defendant will end up with the same pair of potential outcomes $(Y_{1}(z_{c}),Y_{0}(z_{c}))$ as long as he or she is assigned to judges with the same observed characteristics $Z_{c}=z_{c}$. Finally, when the only instrument variable we observe in the data is the identity of the judge $Z_{I}$, then the partial exclusion assumption is equivalent to the original exclusion Assumption 2.2.
The partial monotonicity Assumption 5.3 was initially introduced in Mogstad, Torgovitsky, and Walters (2019). It significantly weakens Assumption 2.3 since it does not require comparing the level of leniency across judges with different observable characteristics. For instance, let $Z_{c}=(Z^{R}_{c},Z^{P}_{c})$ be composed of the following binary variables: $Z^{R}_{c}$ equal to $1$ if the judge is black or Hispanic and 0 if not, while $Z^{P}_{c}$ is $1$ if the judge is from the Republican party and 0 if from the Democratic party. Imposing Assumption 2.3 means it is not possible to have a black democrat judge be more lenient than a white republican judge for some defendants while being less lenient for other defendants, i.e., these two judges may have different cut-off points, but they rank all the defendants in the same order. Mathematically, we can not have both $\mathbb{P}(D(z_{I},1,0)=1,D(z_{I}^{\prime},0,1)=0)>0$ and $\mathbb{P}(D(z_{I}^{\prime},0,1)=1,D(z_{I},1,0)=0)>0$. However, there is a large body of empirical evidence of heterogeneity in the ranking of judges’ leniency across different types of offense or defendants (see Abrams, Bertrand, and Mullainathan, 2012; Stevenson, 2018). This is, however, compatible with the partial monotonicity. Its main advantage is that it no longer requires a uniform ranking of defendants across different judges. Judges’ rankings are allowed to vary with their characteristics $Z_{c}$. Applying the result of Vytlacil (2002), the partial monotonicity condition can be characterized as a partial single threshold-crossing restriction under the independence assumption Assumption 2.1, which we restated below.
@@
Type $Z=(Z_{I},Z_{c})$’s judge treatment assignment mechanism is governed by the following threshold crossing model $D_{z}=1\{\nu(Z_{I},Z_{c})\geq U_{Z_{c}}\}$ for a measurable function $\nu$, where the distribution of $U_{z_{c}}$ is absolutely continuous for all $z_{c}\in\mathcal{Z}_{c}$.
Under Assumptions 2.1 and 5.4, we can apply the standard normalization,
$D(z_{I},z_{c})=1\left\{F_{U_{z_{c}}|Z_{c}}(\nu(z_{I},z_{c})|z_{c})\geq F_{U_{z_{c}}|Z_{c}}(U_{z_{c}}|z_{c})\right\}\equiv 1\left\{P(z_{I},z_{c})\geq V_{z_{c}}\right\},$
@@
The key difference between the STC and the Partial STC is even though $V_{z_{c}}$ follows $Uniform[0,1]$ distribution, each defendant does not face a single $V$. Instead, he or she faces a collection of $\{V_{z_{c}},z_{c}\in\mathcal{Z}_{c}\}$. This unobserved latent variable is now different for judges with distinct observable characteristics. The partial STC has a natural interpretation as an extension of the Roy model (Canay, Mogstad, and Mountjoy, 2024). We can interpret $P(z_{I},z_{c})$ as the perceived gain of incarcerating a defendant by a type $z=(z_{I},z_{c})$ judge, and $V_{z_{c}}$ as the expected cost (but unobserved to the econometrician) of incarcerating a defendant. The particularity of the partial STC is that the expected cost can vary across judges with distinct observable characteristics $z_{c}$, but is fixed within judges with the same $z_{c}$. In the standard monotonicity assumption, the cost $V$ would be the same regardless of the characteristics $(z_{I},z_{c})$. For the same reason, the partial STC is also more reasonable in settings where decision-makers (judges) differ in their preferences and skills (Chan, Gentzkow, and Yu, 2022).
Here, we provide an example of eight judges deciding whether to incarcerate a given defendant to elucidate further the richer heterogeneity enabled by the partial monotonicity (or, equivalently, the partial STC) assumption. We consider the two observable characteristics of the judges introduced earlier, $Z_{c}\equiv(Z_{c}^{R},Z_{c}^{P})\in\{0,1\}\times\{0,1\}$. These two binary observable characteristics result in four types of judges. The eight judges are evenly allocated across these four types.
@@
The left rectangle of Figure 6 shows the benefit and the expected cost of incarcerating the defendant in a separate unit segment for each judge. For example, $p_{11}$ and $p_{11}^{\prime}$ are the benefits of the two black democratic judges with type $Z_{c}=(1,1)$ to incarcerate the defendant. The right rectangle of Figure 6 plots the benefit numbers of all eight judges on the same unit segment. Similarly, $U_{11}$ represents the expected cost of incarcerating the defendant by a black democratic judge: they share the same expected cost or skills. A judge incarcerates the defendant when the corresponding benefit is higher than the expected cost of incarceration. In Figure 6, the judges who incarcerate the defendant are blue-colored, while those who release the defendant are red-colored.
The behavior of the eight judges does not violate Assumption 5.3 or Assumption 5.4. However, the standard monotonicity Assumption 2.3 is clearly violated (right rectangle of Figure 6). Indeed, the judge with propensity score $p_{11}^{\prime}$ incarcerates the defendant (blue-colored), whereas judges with higher propensity scores $p_{01}$, $p_{10}$, or $p_{00}$ do not incarcerate the defendant (red-colored). Note that Assumption 2.3 would not be violated for this group of judges only under one of these two conditions: (i) all four $V_{z_{c}}$ are greater than the maximum of the eight propensities or smaller than the minimum of all eight properties. In other words, when all judges make the same decision regarding this defendant, or (ii) judges who do not incarcerate the defendant must have lower benefit scores than judges who incarcerate the defendant. Moreover, one of these two conditions must hold for all defendants when we impose Assumption 2.3 (or Assumption 2.4).
@@
Assumptions 2.1, 5.2 and 5.4 are weaker than Assumptions 2.1, 2.2 and 2.3. We show that under these weaker conditions, it is still possible to identify meaningful treatment effect parameters.
Theorem 5.3 (Identification under Partial exclusion and monotonicity)
If Assumptions 2.1, 5.2 and 5.4 hold, then:
@@
    (Identification of the LATE). Let $\mathcal{P}_{z_{c}}$ be the support of $P(Z_{I},Z_{c})$ conditioning on $Z_{c}=z_{c}$. Then for any pair $(p,p^{\prime})\in\mathcal{P}_{z_{c}}\times\mathcal{P}_{z_{c}}$ such that $p<p^{\prime}$ we have the following identification results:
    $\frac{\mathbb{E}[g(Y)|P=p^{\prime},Z_{c}=z_{c}]-\mathbb{E}[g(Y)|P=p,Z_{c}=z_{c}]}{p^{\prime}-p}\\ =\mathbb{E}[g(Y_{1}(z_{c}))-g(Y_{0}(z_{c}))|1\{p<V_{z_{c}}\leq p^{\prime}\}].$
@@
    (Identification of the MTE). For any $p\in\mathcal{P}_{z_{c}}$ such that $\mathbb{E}[g(Y)|P=\cdot,Z_{c}=z_{c}]$ is continuously differentiable in the neighborhood of $p$, then,
    $\frac{\partial\mathbb{E}[g(Y)|P=t,Z_{c}=z_{c}]}{\partial t}\big|_{t=p}=\mathbb{E}[g(Y_{1}(z_{c}))-g(Y_{0}(z_{c}))|V_{z_{c}}=p].$
@@
    (Testable restrictions). For any fixed $z_{c}\in\mathcal{Z}_{c}$, $\mathbb{P}(y<Y\leq y^{\prime},D=1|P=p,Z_{c}=z_{c})$ and $\mathbb{P}(y<Y\leq y^{\prime},D=0|P=p,Z_{c}=z_{c})$ are non-decreasing in $p$ for all $p\in\mathcal{P}_{Z_{c}}$ and any $y,y^{\prime}\in\mathcal{Y}$.
The proof of Theorem 5.3 is similar to Theorem 2.1 after conditioning on $Z_{c}=z_{c}$ and therefore omitted. The identification results stated in Theorem 5.3 (i)-(ii) demonstrate that whenever there are two judges with distinct $Z_{I}$ but share the same observed characteristics $Z_{c}=z_{c}$, the conditional Wald estimand identifies the LATE provided the propensity scores for these two judges are different. Moreover, when the distribution of $Z_{I}|Z_{c}=z_{c}$ allows one to take the derivative of $\mathbb{E}[g(Y)|P=\cdot,Z_{c}=z_{c}]$, the conditional LIV estimand identifies the MTE. This identification result is a local version of the standard LATE and MTE identification.
@@
which is a total of fourteen inequalities. However, when invoking our weaker set of assumptions, we have only eight inequalities that characterize the sharp testable implications:
@@
The comparison of the testable implications in Theorems 2.1 and 5.3 confirms that the judge leniency design is more stringent than the conditional judge leniency design. Hence, whenever the standard judge leniency design is rejected, the researcher may rely on its relaxed versions as long as the testable implications derived in Theorem 5.3 are satisfied.